# ShapeMoiré: Channel-Wise Shape-Guided Network for Image Demoiréing

### v1

@@ III-A ShapeConv
Given an input image with a moiré pattern, $I_{moir\acute{e}}$, image demoiréing aims to remove these unwanted patterns and thereby restore the original pristine image. Existing methods [11, 7, 6, 10] often involve an encoder-decoder framework, wherein the conventional convolutional layer serves as the core building block. A convolutional layer takes as input a feature map, either raw pixels or intermediate image features, in a sliding window manner. It typically comprises multiple convolution kernels, which are applied to a patch of images at each window position. To facilitate our discussion, we conceptualize one image patch as a 3D tensor $\mathbb{P}\in R^{K_{h}\times K_{w}\times C_{in}}$, $K_{h}$ and $K_{w}$ are the spatial dimensions of the kernel and $C_{in}$ represents the channel numbers in the input feature map. The output features from the vanilla convolution layer are obtained by,
$\mathbb{F}=Conv(\mathbb{K},\mathbb{P}),$ (1)
@@
$\begin{cases}\mathbb{F}=\mathbb{K}\circ\mathbb{P}\\ \mathbb{F}_{c_{out}}=\sum_{i}^{K_{h}\times K_{w}\times C_{in}}(\mathbb{K}_{i,c_{out}}\times\mathbb{P}_{i}).\end{cases}$ (2)
@@
$\mathbb{P}_{B}=\overline{m}(\mathbb{P}),$ (3)
@@
$\mathbb{F}$ $=ShapeConv(\mathbb{K},\mathbb{W}_{B},\mathbb{W}_{S},\mathbb{P})$ (4)
@@
$\begin{cases}\textbf{P}_{\textbf{B}}=\mathbb{W}_{B}\diamond\mathbb{P}_{B}\\ \textbf{P}_{\textbf{B}_{1,1,c_{in}}}=\mathbb{W}_{B}\times\mathbb{P}_{B_{1,1,c_{in}}},\end{cases}$ (5)
$\begin{cases}\textbf{P}_{\textbf{S}}=\mathbb{W}_{S}\ast\mathbb{P}_{S}\\ \textbf{P}_{\textbf{S}_{{k_{h}},{k_{w}},{c_{in}}}}=\sum_{i}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{i,{k_{h}},{k_{w}},{c_{in}}}}\times\mathbb{P}_{S_{i,{c_{in}}}}),\end{cases}$ (6)
@@
III-B ShapeConv in Training and Inference
@@
$\mathbb{F}$ $=ShapeConv(\mathbb{K},\mathbb{W}_{B},\mathbb{W}_{S},\mathbb{P})$ (7)
@@
$\mathbb{F}$ $=ShapeConv(\mathbb{K},\mathbb{W}_{B},\mathbb{W}_{S},\mathbb{P})$ (8)
@@
$\mathbb{F}_{c_{out}}$ $=\sum_{i}^{K_{h}\times K_{w}\times C_{in}}(\mathbb{K}_{i,{c_{out}}}\times\textbf{P}_{\textbf{BS}_{i}})$ (9)
@@
Figure 3: Comparison of vanilla convolution layer and ShapeConv within a patch $\mathbb{P}$. In this figure, $K_{h}=K_{w}=2$, $C_{in}=3$, and $C_{out}=2$, “+” denotes element-wise addition. (a) Vanilla convolution with kernel $\mathbb{K}$; (b) ShapeConv with folding the $\mathbb{W}_{B}$ and $\mathbb{W}_{S}$ into $\textbf{K}_{\textbf{BS}}$; (c) The computation of $\textbf{K}_{\textbf{BS}}$ from $\mathbb{K}$, $\mathbb{W}_{B}$ and $\mathbb{W}_{S}$.
@@
TABLE I: Performance comparison with state-of-the-art methods on LCDMoiré, TIP2018, FHDMi, and UHDM datasets. ($\uparrow$) and ($\downarrow$) indicate better performance with larger and smaller numbers, respectively. Red: best and Blue: second-best.
@@
Dataset Metrics Input  [10]  [5]  [11]  [12]  [7]  [7]  [32]  [33] ShapeMoiré
@@
III-C Shape-Architecture
Method Intuition
[…]
$\mathcal{L}_{base}=L_{1}(I_{output},I_{gt})+\lambda L_{p}(I_{output},I_{gt}),$ (10)
where $\lambda$ is a hyperparameter that balances between $L_{1}$ and $L_{p}$.
Shape-Architecture
$I^{Shape}_{moir\acute{e}}$ $=\mathcal{M}(I_{moir\acute{e}}),$ (11)
$I^{Shape}_{gt}$ $=\mathcal{M}(I_{gt}),$
$\mathcal{L}_{shape}=L_{1}(I^{Shape}_{output},I^{Shape}_{gt})+\lambda L_{p}(I^{Shape}_{output},I^{Shape}_{gt}).$ (12)
The final loss function $\mathcal{L}_{total}$ is:
$\mathcal{L}_{total}=\mathcal{L}_{base}+\mathcal{L}_{shape}.$ (13)
Shape-Architecture During Inference
III-D ShapeMoiré
TABLE II: Performance comparison with different baseline methods on four datasets. An asterisk (*) indicates the performance we reproduced on the respective datasets using the source code provided by different methods, noting slight discrepancies from the original publications.
UHDM [7] FHDMi [6] TIP2018 [10] LCDMoiré [9]
@@ III-D ShapeMoiré
ESDNet [7] $\Delta$ 0.344$\uparrow$ .0033$\uparrow$ .0021$\downarrow$ 0.236$\uparrow$ .0010$\uparrow$ .0030$\downarrow$ 0.071$\uparrow$ .0005$\uparrow$ 0.251$\uparrow$ .0004$\uparrow$ 0
@@
ESDNet-L [7] $\Delta$ 0.394$\uparrow$ .0024$\uparrow$ .0053$\downarrow$ 0.256$\uparrow$ .0039$\uparrow$ .0015$\downarrow$ 0.065$\uparrow$ .0006$\uparrow$ 1.014$\uparrow$ .0005$\uparrow$ 0
@@
WDNet [11] $\Delta$ 0.701$\uparrow$ .1025$\uparrow$ .0999$\downarrow$ 1.021$\uparrow$ .0187$\uparrow$ .0277$\downarrow$ 0.500$\uparrow$ .0154$\uparrow$ 1.084$\uparrow$ .0061$\uparrow$ 0
Baseline* 17.812 .7294 .5561 19.313 .7375 .3395 24.518 .8073 29.321 .9482 1.426
ShapeMoiré 18.036 .7336 .5185 19.615 .7495 .2889 25.381 .8234 29.649 .9570 1.426
DMCNN [10] $\Delta$ 0.224$\uparrow$ .0042$\uparrow$ .0376$\downarrow$ 0.302$\uparrow$ .0120$\uparrow$ .0506$\downarrow$ 0.863$\uparrow$ .0161$\uparrow$ 0.328$\uparrow$ .0088$\uparrow$ 0
IV Experiments
(1) LCDMoiré Dataset [9]. As a part of the AIM19 image demoiréing challenge, the LCDMoiré dataset comprises 10,200 synthetically generated image pairs. It is structured to include 10,000 images for training purposes, 100 images for validation, and another 100 for testing. Each image is of high resolution, with dimensions of 1024 $\times$ 1024 pixels.
(2) TIP2018 Dataset [10]. This dataset consists of real photographs obtained by capturing images from ImageNet [35] displayed on computer screens, ensuring each image has a resolution of 256 $\times$ 256 pixels.
(3) FHDMi Dataset [6]. It includes a substantial collection of 9,981 training image pairs and 2,019 testing pairs. All images in this dataset are in full high-definition resolution, measuring 1920 $\times$ 1080 pixels.
(4) UHDM Dataset [7]. It stands out for its 4K resolution demoiréing capability. It is an ultra-high-resolution dataset featuring 5,000 image pairs that span a broad spectrum of scenes, including landscapes, sports, video clips, and documents. The dataset is notable for its diverse moiré patterns, produced using various device combinations and viewpoints. Each image in this dataset is 3840 $\times$ 2160 pixels.
Evaluation Protocols. For quantitative evaluation, we employed widely recognized metrics, including PSNR (Peak Signal-to-Noise Ratio), SSIM (Structure Similarity) [36], and LPIPS (Learned Perceptual Image Patch Similarity) [37]. In addition to these performance-related metrics, we also considered the network’s parameter count, denoted as Param.(M), which indicates the number of network parameters in millions. This metric is pertinent to the overall memory usage of the network. It is important to note that previous methods [10, 16, 5, 11, 7] have exclusively reported PSNR and SSIM metrics on the TIP2018 and LCDMoiré datasets. In line with this convention, we adhere to the same setup for our comparisons.
@@ IV Experiments
Implementation Details. We adopted several popular approaches [10, 11, 7] as our baseline to demonstrate both the effectiveness and generalization capability of our proposed method. We strictly followed the original implementations of these baseline methods, except for the model components of ShapeMoiré. This guarantees that the observed performance improvements are solely attributable to the application of ShapeMoiré, yet not influenced by other factors. Unless otherwise noted, the baseline model is ESDNet-L [7]. Detailed implementations of our method can be found in our released code.
Figure 4: Visualization results from the UHDM dataset. Larger boxes represent amplified views of the corresponding smaller colored border boxes, offering a clearer perspective when enlarged. For samples A and B, the top and bottom rows respectively depict the visualization results using WDNet [11] and ESDNet-L [7] as baseline methods.
IV-A Comparison with State-of-the-Art Methods
@@
We compared our method with state-of-the-art techniques and reported the results in Table I. As evident from the table, our ShapeMoiré significantly outperforms existing methods across most scenarios. It is noteworthy that our approach establishes a new SOTA across all four datasets in terms of the PSNR metric, which quantifies the level of noise or distortion compared to the original undistorted image. The results highlight that the modeled shape features in our Shapemoiré play a crucial role in identifying the causal factors essential for image demoiréing. Furthermore, ShapeMoiré demonstrates superior performance compared to existing approaches on the UHDM dataset, emphasizing the effectiveness of our proposed method for Ultra-High-Definition images.
IV-B Generalization to Diverse Architectures
@@
The proposed ShapeMoiré is a versatile approach that can be seamlessly integrated into existing methods for image demoiréing. To test its generalization capability, we applied our ShapeMoiré to four representative demoiréing methods, namely ECDNet [7], ECDNet-L [7], WDnet [11], and DMCNN [10]. Table II demonstrates that ShapeMoiré consistently delivers significant performance improvements across ALL settings. Moreover, our method introduces no additional inference cost compared to baselines, i.e., the added parameters are zero. It is worth noting that, as observed in Table II, the performance of current methods mostly correlates positively with the number of parameters. However, ShapeMoiré demonstrates the ability to enhance performance without increasing the network’s parameter count. For instance, our method achieves comparable (or even higher PSNR) performance to ESDNet-L (which has almost twice the parameter count of ESDNet) on the UHDM dataset, even when keeping the parameter count consistent with ESDNet.
IV-C Ablation Study

### v2

@@ 3.1. ShapeConv
Given an input image with a moiré pattern, $I_{moir\acute{e}}$, image demoiréing aims to remove these unwanted patterns and thereby restore the original pristine image. Existing methods (Liu et al., 2020; Yu et al., 2022; He et al., 2020; Sun et al., 2018) often involve an encoder-decoder framework, wherein the conventional convolutional layer serves as the core building block. A convolutional layer takes as input a feature map, either raw pixels or intermediate image features, in a sliding window manner. It typically comprises multiple convolution kernels, which are applied to a patch of images at each window position. To facilitate our discussion, we conceptualize one image patch as a 3D tensor $\mathbb{P}\in R^{K_{h}\times K_{w}\times C_{in}}$, $K_{h}$ and $K_{w}$ are the spatial dimensions of the kernel and $C_{in}$ represents the channel numbers in the input feature map. The output features from the vanilla convolution layer are obtained by,
(1) $\mathbb{F}=Conv(\mathbb{K},\mathbb{P}),$
@@
(2) $\begin{cases}\mathbb{F}=\mathbb{K}\circ\mathbb{P}\\ \mathbb{F}_{c_{out}}=\sum_{i}^{K_{h}\times K_{w}\times C_{in}}(\mathbb{K}_{i,c_{out}}\times\mathbb{P}_{i}).\end{cases}$
@@
(3) $\mathbb{P}_{B}=\overline{m}(\mathbb{P}),$
@@
(4) $\mathbb{F}$ $=ShapeConv(\mathbb{K},\mathbb{W}_{B},\mathbb{W}_{S},\mathbb{P})$
@@
(5) $\begin{cases}\textbf{P}_{\textbf{B}}=\mathbb{W}_{B}\diamond\mathbb{P}_{B}\\ \textbf{P}_{\textbf{B}_{1,1,c_{in}}}=\mathbb{W}_{B}\times\mathbb{P}_{B_{1,1,c_{in}}},\end{cases}$
(6) $\begin{cases}\textbf{P}_{\textbf{S}}=\mathbb{W}_{S}\ast\mathbb{P}_{S}\\ \textbf{P}_{\textbf{S}_{{k_{h}},{k_{w}},{c_{in}}}}=\sum_{i}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{i,{k_{h}},{k_{w}},{c_{in}}}}\times\mathbb{P}_{S_{i,{c_{in}}}}),\end{cases}$
@@
3.2. ShapeConv in Training and Inference
@@
(7) $\mathbb{F}$ $=ShapeConv(\mathbb{K},\mathbb{W}_{B},\mathbb{W}_{S},\mathbb{P})$
@@
(8) $\mathbb{F}$ $=ShapeConv(\mathbb{K},\mathbb{W}_{B},\mathbb{W}_{S},\mathbb{P})$
@@
(9) $\mathbb{F}_{c_{out}}$ $=\sum_{i}^{K_{h}\times K_{w}\times C_{in}}(\mathbb{K}_{i,{c_{out}}}\times\textbf{P}_{\textbf{BS}_{i}})$
@@
Figure 3. Comparison of vanilla convolution layer and ShapeConv within a patch $\mathbb{P}$. In this figure, $K_{h}=K_{w}=2$, $C_{in}=3$, and $C_{out}=2$, “+” denotes element-wise addition. (a) Vanilla convolution with kernel $\mathbb{K}$; (b) ShapeConv with folding the $\mathbb{W}_{B}$ and $\mathbb{W}_{S}$ into $\textbf{K}_{\textbf{BS}}$; (c) The computation of $\textbf{K}_{\textbf{BS}}$ from $\mathbb{K}$, $\mathbb{W}_{B}$ and $\mathbb{W}_{S}$.
@@
3.3. Shape-Architecture
Method Intuition
[…]
(10) $\mathcal{L}_{base}=L_{1}(I_{output},I_{gt})+\lambda L_{p}(I_{output},I_{gt}),$
where $\lambda$ is a hyperparameter that balances between $L_{1}$ and $L_{p}$.
Shape-Architecture
(11) $I^{Shape}_{moir\acute{e}}$ $=\mathcal{M}(I_{moir\acute{e}}),$
$I^{Shape}_{gt}$ $=\mathcal{M}(I_{gt}),$
(12) $\mathcal{L}_{shape}=L_{1}(I^{Shape}_{output},I^{Shape}_{gt})+\lambda L_{p}(I^{Shape}_{output},I^{Shape}_{gt}).$
The final loss function $\mathcal{L}_{total}$ is:
(13) $\mathcal{L}_{total}=\alpha\mathcal{L}_{base}+\beta\mathcal{L}_{shape},$
where $\alpha$ and $\beta$ are the weighting coefficients for the loss terms $\mathcal{L}_{base}$ and $\mathcal{L}_{shape}$, respectively.
Shape-Architecture During Inference
Relationship between ShapeConv and Shape-Architecture.
While ShapeConv and Shape-Architecture operate at different levels—patch-level and image-level respectively—they are intrinsically aligned through the shared notion of shape decomposition. Specifically, both components rely on the same shape operator $\mathcal{M}$ to extract the relative information from input data by subtracting the channel-wise mean. ShapeConv uses this operator within local patches to enhance feature discrimination, whereas Shape-Architecture applies it globally to form a parallel shape stream at the image level. This design ensures consistency in how shape information is perceived and utilized throughout the network. Together, they form a hierarchical shape-aware framework, where ShapeConv captures local fine-grained shape variations and Shape-Architecture reinforces the model’s global structural awareness. This joint modeling helps the network better suppress moiré patterns that exhibit both local distortions and large-scale repetitive interference.
3.4. ShapeMoiré
Table 1. Performance comparison with state-of-the-art methods on LCDMoiré, TIP2018, FHDMi, and UHDM datasets. ($\uparrow$) and ($\downarrow$) indicate better performance with larger and smaller numbers, respectively. Red: best and Blue: second-best.
@@ 3.4. ShapeMoiré
Dataset Metrics Input  (Sun et al., 2018)  (He et al., 2019)  (Liu et al., 2020)  (Zheng et al., 2020)  (Yu et al., 2022)  (Yu et al., 2022)  (Zhang et al., 2023)  (Xu et al., 2023b) ShapeMoiré
@@
4. Experiments
(1) LCDMoiré Dataset (Yuan et al., 2019). As a part of the AIM19 image demoiréing challenge, the LCDMoiré dataset comprises 10,200 synthetically generated image pairs. It is structured to include 10,000 images for training purposes, 100 images for validation, and another 100 for testing. Each image is of high resolution, with dimensions of 1024 $\times$ 1024 pixels.
(2) TIP2018 Dataset (Sun et al., 2018). This dataset consists of real photographs obtained by capturing images from ImageNet (Deng et al., 2009) displayed on computer screens, ensuring each image has a resolution of 256 $\times$ 256 pixels.
(3) FHDMi Dataset (He et al., 2020). It includes a substantial collection of 9,981 training image pairs and 2,019 testing pairs. All images in this dataset are in full high-definition resolution, measuring 1920 $\times$ 1080 pixels.
(4) UHDM Dataset (Yu et al., 2022). It stands out for its 4K resolution demoiréing capability. It is an ultra-high-resolution dataset featuring 5,000 image pairs that span a broad spectrum of scenes, including landscapes, sports, video clips, and documents. The dataset is notable for its diverse moiré patterns, produced using various device combinations and viewpoints. Each image in this dataset is 3840 $\times$ 2160 pixels.
Evaluation Protocols. For quantitative evaluation, we employed widely recognized metrics, including PSNR (Peak Signal-to-Noise Ratio), SSIM (Structure Similarity) (Wang et al., 2004), and LPIPS (Learned Perceptual Image Patch Similarity) (Zhang et al., 2018a). In addition to these performance-related metrics, we also considered the network’s parameter count, denoted as Param.(M), which indicates the number of network parameters in millions. This metric is pertinent to the overall memory usage of the network. It is important to note that previous methods (Sun et al., 2018; Cheng et al., 2019; He et al., 2019; Liu et al., 2020; Yu et al., 2022) have exclusively reported PSNR and SSIM metrics on the TIP2018 and LCDMoiré datasets. In line with this convention, we adhere to the same setup for our comparisons.
Table 2. Performance comparison with different baseline methods on four datasets. An asterisk (*) indicates the performance we reproduced on the respective datasets using the source code provided by different methods, noting slight discrepancies from the original publications.
UHDM (Yu et al., 2022) FHDMi (He et al., 2020) TIP2018 (Sun et al., 2018) LCDMoiré (Yuan et al., 2019)
@@ 4. Experiments
ESDNet (Yu et al., 2022) $\Delta$ 0.344$\uparrow$ .0033$\uparrow$ .0021$\downarrow$ 0.236$\uparrow$ .0010$\uparrow$ .0030$\downarrow$ 0.071$\uparrow$ .0005$\uparrow$ 0.251$\uparrow$ .0004$\uparrow$ 0
@@
ESDNet-L (Yu et al., 2022) $\Delta$ 0.394$\uparrow$ .0024$\uparrow$ .0053$\downarrow$ 0.256$\uparrow$ .0039$\uparrow$ .0015$\downarrow$ 0.065$\uparrow$ .0006$\uparrow$ 1.014$\uparrow$ .0005$\uparrow$ 0
@@
WDNet (Liu et al., 2020) $\Delta$ 0.701$\uparrow$ .1025$\uparrow$ .0999$\downarrow$ 1.021$\uparrow$ .0187$\uparrow$ .0277$\downarrow$ 0.500$\uparrow$ .0154$\uparrow$ 1.084$\uparrow$ .0061$\uparrow$ 0
Baseline* 17.812 .7294 .5561 19.313 .7375 .3395 24.518 .8073 31.425 .9552 1.426
ShapeMoiré 18.036 .7336 .5185 19.615 .7495 .2889 25.381 .8234 31.789 .9562 1.426
DMCNN (Sun et al., 2018) $\Delta$ 0.224$\uparrow$ .0042$\uparrow$ .0376$\downarrow$ 0.302$\uparrow$ .0120$\uparrow$ .0506$\downarrow$ 0.863$\uparrow$ .0161$\uparrow$ 0.364$\uparrow$ .0010$\uparrow$ 0
@@
Implementation Details. We adopted several popular approaches (Sun et al., 2018; Liu et al., 2020; Yu et al., 2022) as our baseline to demonstrate both the effectiveness and generalization capability of our proposed method. We strictly followed the original implementations of these baseline methods, except for the model components of ShapeMoiré. This guarantees that the observed performance improvements are solely attributable to the application of ShapeMoiré, yet not influenced by other factors. Unless otherwise noted, the baseline model is ESDNet-L (Yu et al., 2022). Detailed implementations of our method can be found in our released code.
4.1. Comparison with State-of-the-Art Methods
@@
We compared our method with state-of-the-art techniques and reported the results in Table 1. As evident from the table, our ShapeMoiré significantly outperforms existing methods across most scenarios. It is noteworthy that our approach establishes a new SOTA across all four datasets in terms of the PSNR metric, which quantifies the level of noise or distortion compared to the original undistorted image. The results highlight that the modeled shape features in our Shapemoiré play a crucial role in identifying the causal factors essential for image demoiréing. Furthermore, ShapeMoiré demonstrates superior performance compared to existing approaches on the UHDM dataset, emphasizing the effectiveness of our proposed method for Ultra-High-Definition images.
4.2. Generalization to Diverse Architectures
@@
The proposed ShapeMoiré is a versatile approach that can be seamlessly integrated into existing methods for image demoiréing. To test its generalization capability, we applied our ShapeMoiré to four representative demoiréing methods, namely ECDNet (Yu et al., 2022), ECDNet-L (Yu et al., 2022), WDnet (Liu et al., 2020), and DMCNN (Sun et al., 2018). Table 2 demonstrates that ShapeMoiré consistently delivers significant performance improvements across ALL settings. Moreover, our method introduces no additional inference cost compared to baselines, i.e., the added parameters are zero. It is worth noting that, as observed in Table 2, the performance of current methods mostly correlates positively with the number of parameters. However, ShapeMoiré demonstrates the ability to enhance performance without increasing the network’s parameter count. For instance, our method achieves comparable (or even higher PSNR) performance to ESDNet-L (which has almost twice the parameter count of ESDNet) on the UHDM dataset, even when keeping the parameter count consistent with ESDNet.
4.3. Generalization to Image Deblurring
Table 3. Dual-pixel Defocus Deblurring comparisons on the DPDD Dataset (Abuolaim and Brown, 2020). The test set of DPDD contains 37 indoor scenes and 39 outdoor scenes.
Archi- Method Indoor Scenes Outdoor Scenes Combined
[…]
MIRNet-v2 (Zamir et al., 2022b) Baseline 28.95 .8800 .1585 23.66 .7549 .2109 26.23 .8158 .1854
ShapreMoiré 29.05 .8841 .1507 23.77 .7569 .1974 26.34 .8188 .1747
$\Delta$ 0.10$\uparrow$ .0041$\uparrow$ .0078$\downarrow$ 0.11$\uparrow$ .020$\uparrow$ .0135$\downarrow$ 0.11$\uparrow$ .0030$\uparrow$ .0107$\downarrow$
To further evaluate the generalization ability of the proposed method beyond the demoiréing task, we extend our experiments to the dual-pixel defocus deblurring task using the DPDD dataset (Abuolaim and Brown, 2020). We adopt MIRNet-v2 (Zamir et al., 2022b) as the baseline architecture and integrate our proposed ShapeConv and Shape-Architecture modules. As presented in Table 3, the ShapeMoiré-enhanced model consistently outperforms the baseline across all three metrics—PSNR, SSIM, and LPIPS—on both indoor and outdoor subsets. Specifically, on the combined test set, our method improves PSNR from 26.23 to 26.34, SSIM from .8158 to .8188, and reduces LPIPS from .1854 to .1747. These improvements indicate not only better pixel-level fidelity but also enhanced perceptual quality.
The consistent performance gains across different domains and evaluation metrics suggest that ShapeMoiré is a versatile and effective enhancement module that can be generalized to broader image restoration tasks beyond demoiréing.
4.4. Ablation Study